OpenStudy (anonymous):

Which set of three numbers could represent the lengths of the sides of a right triangle? A. 7, 9, 11 B. 8, 15, 17 C. 10, 15, 20 D. 12, 18, 22

OpenStudy (anonymous):

I tried all but only two remain A and D and Im confused

OpenStudy (anonymous):

B. 8, 15, 17 @TakeA.I.M.

OpenStudy (anonymous):

how

OpenStudy (anonymous):

Can you explain @dpasingh

OpenStudy (anonymous):

\[17^2= 289\] \[8^2+15^2= 64+225 = 289\] Therefore \[8^2+15^2=17^2\] Thus 8, 15, 17 are the lengths of the sides of a right triangle.

OpenStudy (anonymous):

@TakeA.I.M.

OpenStudy (anonymous):

Ok thnx